# Global Convergence for Average Reward Constrained MDPs with Primal-Dual Actor Critic Algorithm

### v1

@@ 5 Conclusion
In this paper, we investigate the infinite-horizon average reward Constrained Markov Decision Processes (CMDPs) with general policy parametrization. We propose a novel algorithm, the “Primal-dual natural actor-critic," which efficiently manages constraints while achieving a global convergence rate of $\tilde{\mathcal{O}}(1/\sqrt{T})$, aligning with the theoretical lower bound for Markov Decision Processes (MDPs). We also provide results when the mixing time is unknown. Future work includes tightening the gap when the mixing time is unknown.
@@ Appendix B Preliminary Results for Global Convergence Analysis
where $R_{0}=\lambda_{P}^{-3}\Lambda_{q}^{2}\sigma_{P}^{2}+\lambda_{P}^{-1}\sigma_{q}^{2}$, $R_{1}=\lambda_{P}^{-2}\big{[}\delta_{P}^{2}\lambda_{P}^{-2}\Lambda_{q}^{2}+\delta_{q}^{2}\big{]}$, and $\tilde{\operatorname*{\mathcal{O}}}(\cdot)$ hides logarithmic factors of $H$. Moreover,
@@
-$\langle$- ⟨ $\mathbb{E}[x_{h}]-x^{*},\mathbb{E}_{h}\left[\mathbb{E}[g_{h}]-P(\mathbb{E}[x_{h}]-x^{*})\right]\rangle$
@@
=$\overset{(b)}{=}\alpha[\mathcal{L}(\pi^{*},\lambda_{k})-\mathcal{L}(\theta_{k},\lambda_{k})]+\alpha\mathbb{E}_{(s,a)\sim\nu^{\pi^{*}}}\bigg{[}\nabla_{\theta}\log\pi_{\theta_{k}}(a|s)\cdot\omega^{*}_{k}-A_{r+\lambda_{k}c}^{\pi_{\theta_{k}}}(s,a)\bigg{]}$
@@
where $\mathcal{L}(\pi^{*},\lambda_{k})\coloneqq J_{r}^{\pi^{*}}+\lambda_{k}J_{c}^{\pi^{*}}$, the relations $(a)$, $(b)$ result from Assumption 4.4$(b)$ and Lemma C.1, respectively. Inequality $(c)$ arises from the convexity of the function $f(x)=x^{2}$. Lastly, $(d)$ is a consequence of Assumption 4.3. By taking expectations on both sides, we derive:
@@
where $(a)$ follows from Assumption 4.4$(a)$. Rearranging the terms, we get,
@@
Adding the above inequality from $k=0$ to $K-1$, using the non-negativity of KL divergence and dividing the resulting expression by $K$, we obtain the final result.
@@
To prove the other statements, recall the definition of ^$\hat{\nabla}_{\theta}J_{g}(\theta_{k},\xi_{g}^{k};\cdot)$ from (22). Note the following relations for arbitrary $\theta_{k},\xi_{g}^{k}=[\eta_{g}^{k},\zeta_{g}^{k}]$.
$\mathbb{E}_{\theta_{k}}$ [^$\left[\hat{\nabla}_{\theta}J_{g}(\theta_{k},\xi_{g}^{k};z)\right]-\nabla_{\theta}J_{g}(\theta_{k})$
@@
=$\overset{(a)}{=}\underbrace{\mathbb{E}_{\theta_{k}}\left[\bigg{\{}\eta_{g}^{*}(\theta_{k})-\eta^{k}_{g}+\langle\phi_{g}(s^{\prime})-\phi_{g}(s),\zeta^{k}_{g}-\zeta_{g}^{*}(\theta_{k})\rangle\bigg{\}}\nabla_{\theta}\log{\pi_{\theta_{k}}}(a|s)\right]}_{T_{0}}$
@@
$\left\lVert T_{0}\right\rVert^{2}$ =$\overset{(a)}{=}\operatorname*{\mathcal{O}}\left(G_{1}^{2}\left\lVert\xi^{k}_{g}-\xi_{g}^{*}(\theta_{k})\right\rVert^{2}\right),~{}\left\lVert T_{1}\right\rVert^{2}\overset{(b)}{=}\operatorname*{\mathcal{O}}\left(G_{1}^{2}\epsilon_{\mathrm{app}}\right),~{}\text{and}~{}T_{2}\overset{(c)}{=}0$ (60)
@@
^$\left\lVert\hat{\nabla}_{\theta}J_{g}(\theta_{k},\xi_{g}^{k};z^{j}_{kh})-\mathbb{E}_{\theta_{k}}\left[\hat{\nabla}_{\theta}J_{g}(\theta_{k},\xi_{g}^{k};z)\right]\right\rVert^{2}\overset{(a)}{\leq}\sigma_{k,g}^{2}=\operatorname*{\mathcal{O}}\left(G_{1}^{2}\left\lVert\xi^{k}_{g}\right\rVert^{2}\right)$ (62)
@@
=$\overset{(a)}{=}\underbrace{\mathbb{E}_{\theta_{k}}\left[\bigg{\{}\eta_{g}^{*}(\theta_{k})-\mathbb{E}_{k}[\eta_{g}^{k}]+\langle\phi_{g}(s^{\prime})-\phi_{g}(s),\mathbb{E}_{k}[\zeta_{g}^{k}]-\zeta^{*}_{g}(\theta_{k})\rangle\bigg{\}}\nabla_{\theta}\log{\pi_{\theta_{k}}}(a|s)\right]}_{T_{0}}+$
@@
$\left\lVert T_{0}\right\rVert^{2}$ =$\overset{(a)}{=}\operatorname*{\mathcal{O}}\left(G_{1}^{2}\left\lVert\mathbb{E}_{k}[\xi_{g}^{k}]-\xi_{g}^{*}(\theta_{k})\right\rVert^{2}\right),~{}\left\lVert T_{1}\right\rVert^{2}\overset{(b)}{=}\operatorname*{\mathcal{O}}\left(G_{1}^{2}\epsilon_{\mathrm{app}}\right),~{}\text{and}~{}T_{2}\overset{(c)}{=}0$ (63)
@@ F.0.1 Rate of Convergence of the Objective
$-\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\bigg{[}\lambda_{k}(J_{c}^{\pi^{*}}-J_{c}(\theta_{k}))\bigg{]}\leq\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\bigg{[}\lambda_{k}(J_{c}(\theta_{k})-\eta_{c}^{k})\bigg{]}+\frac{\beta}{K}\sum_{k=0}^{K-1}\mathbb{E}[\eta_{c}^{k}]^{2}$ (74)
@@ F.0.2 Rate of Constraint Violation
$0\leq\frac{1}{K}\left|\lambda_{K}-\frac{2}{\delta}\right|^{2}\leq\frac{1}{K}|{{\lambda_{0}-\bar{\lambda}}}|^{2}-\frac{2\beta}{K}\sum_{k=0}^{K-1}J_{c}(\theta_{k})\left(\lambda_{k}-\frac{2}{\delta}\right)+\beta^{2}$ (78)
@@ Appendix G Some Auxiliary Lemmas for the Proofs
[33, Lemma 14] For any ergodic MDP with mixing time $\tau_{\mathrm{mix}}$, the following holds $\forall(s,a)\in\mathcal{S}\times\mathcal{A}$, any policy $\pi$ and $\forall g\in\{r,c\}$.
@@
Applying Lemma G.1$(b)$ and Assumption 4.4$(a)$, we get
@@
[16, Lemma 3] We restate the problem (2) below for convenience.
@@
Notice that in Eq. (12), the dual update is projected onto the set $[0,\frac{2}{\delta}]$ because the optimal dual variable for the parameterized problem is bounded in Lemma G.4. We note that our dual updating technique remains the same as [10], however we were able to achieve an improvement of global convergence rate due to the use of natural policy gradient and a more prudent choice of stepsizes ($\alpha,\beta$).

### v2

@@ 5 Conclusion
In this paper, we investigate the infinite-horizon average reward Constrained Markov Decision Processes (CMDPs) with general policy parametrization. We propose a novel algorithm, the “Primal-dual natural actor-critic," which efficiently manages constraints while achieving a global convergence rate of $\tilde{\mathcal{O}}(1/\sqrt{T})$, aligning with the theoretical lower bound for Markov Decision Processes (MDPs). We also extend our analysis to the setting with unknown mixing time. Future directions include narrowing the performance gap in this setting, parameterizing the critic using neural networks as in gaur2024closing ; ganesh2025order , and relaxing the ergodicity assumption following ganesh2025regret .
6 Acknowledgment
The work was supported in part by the National Science Foundation under grant CCF-2149588, Office of Naval Research under grant N00014-23-1-2532, and Cisco Systems, Inc.
@@ Appendix B Preliminary Results for Global Convergence Analysis
[…]
@@
-$\langle$ $\mathbb{E}[x_{h}]-x^{*},\mathbb{E}_{h}\left[\mathbb{E}[g_{h}]-P(\mathbb{E}[x_{h}]-x^{*})\right]\rangle$
@@
$\overset{(b)}{=}\alpha[\mathcal{L}(\pi^{*},\lambda_{k})-\mathcal{L}(\theta_{k},\lambda_{k})]+\alpha\mathbb{E}_{(s,a)\sim\nu^{\pi^{*}}}\bigg[\nabla_{\theta}\log\pi_{\theta_{k}}(a|s)\cdot\omega^{*}_{k}-A_{r+\lambda_{k}c}^{\pi_{\theta_{k}}}(s,a)\bigg]$
@@
where $\mathcal{L}(\pi^{*},\lambda_{k})\coloneqq J_{r}^{\pi^{*}}+\lambda_{k}J_{c}^{\pi^{*}}$, the relations $(a)$, $(b)$ result from Assumption 4.4(b)(b) and Lemma C.1, respectively. Inequality $(c)$ arises from the convexity of the function $f(x)=x^{2}$. Lastly, $(d)$ is a consequence of Assumption 4.3. By taking expectations on both sides, we derive:
@@
where $(a)$ follows from Assumption 4.4(a)(a). Rearranging the terms, we get,
@@
Adding the above inequality from $k=0$ to K1K-1, using the non-negativity of KL divergence and dividing the resulting expression by $K$, we obtain the final result.
@@
To prove the other statements, recall the definition of $\hat{\nabla}_{\theta}J_{g}(\theta_{k},\xi_{g}^{k};\cdot)$ from (22). Note the following relations for arbitrary $\theta_{k},\xi_{g}^{k}=[\eta_{g}^{k},\zeta_{g}^{k}]$.
$\mathbb{E}_{\theta_{k}}$ $\left[\hat{\nabla}_{\theta}J_{g}(\theta_{k},\xi_{g}^{k};z)\right]-\nabla_{\theta}J_{g}(\theta_{k})$
@@
$\overset{(a)}{=}\underbrace{\mathbb{E}_{\theta_{k}}\left[\bigg\{\eta_{g}^{*}(\theta_{k})-\eta^{k}_{g}+\langle\phi_{g}(s^{\prime})-\phi_{g}(s),\zeta^{k}_{g}-\zeta_{g}^{*}(\theta_{k})\rangle\bigg\}\nabla_{\theta}\log{\pi_{\theta_{k}}}(a|s)\right]}_{T_{0}}$
@@
$\left\lVert T_{0}\right\rVert^{2}$ $\overset{(a)}{=}\operatorname*{\mathcal{O}}\left(G_{1}^{2}\left\lVert\xi^{k}_{g}-\xi_{g}^{*}(\theta_{k})\right\rVert^{2}\right),~\left\lVert T_{1}\right\rVert^{2}\overset{(b)}{=}\operatorname*{\mathcal{O}}\left(G_{1}^{2}\epsilon_{\mathrm{app}}\right),~\text{and}~T_{2}\overset{(c)}{=}0$ (60)
@@
$\left\lVert\hat{\nabla}_{\theta}J_{g}(\theta_{k},\xi_{g}^{k};z^{j}_{kh})-\mathbb{E}_{\theta_{k}}\left[\hat{\nabla}_{\theta}J_{g}(\theta_{k},\xi_{g}^{k};z)\right]\right\rVert^{2}\overset{(a)}{\leq}\sigma_{k,g}^{2}=\operatorname*{\mathcal{O}}\left(G_{1}^{2}\left\lVert\xi^{k}_{g}\right\rVert^{2}\right)$ (62)
@@
$\overset{(a)}{=}\underbrace{\mathbb{E}_{\theta_{k}}\left[\bigg\{\eta_{g}^{*}(\theta_{k})-\mathbb{E}_{k}[\eta_{g}^{k}]+\langle\phi_{g}(s^{\prime})-\phi_{g}(s),\mathbb{E}_{k}[\zeta_{g}^{k}]-\zeta^{*}_{g}(\theta_{k})\rangle\bigg\}\nabla_{\theta}\log{\pi_{\theta_{k}}}(a|s)\right]}_{T_{0}}+$
@@
$\left\lVert T_{0}\right\rVert^{2}$ $\overset{(a)}{=}\operatorname*{\mathcal{O}}\left(G_{1}^{2}\left\lVert\mathbb{E}_{k}[\xi_{g}^{k}]-\xi_{g}^{*}(\theta_{k})\right\rVert^{2}\right),~\left\lVert T_{1}\right\rVert^{2}\overset{(b)}{=}\operatorname*{\mathcal{O}}\left(G_{1}^{2}\epsilon_{\mathrm{app}}\right),~\text{and}~T_{2}\overset{(c)}{=}0$ (63)
@@ F.0.1 Rate of Convergence of the Objective
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\bigg[\lambda_{k}(J_{c}^{\pi^{*}}-J_{c}(\theta_{k}))\bigg]\leq\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\bigg[\lambda_{k}(J_{c}(\theta_{k})-\eta_{c}^{k})\bigg]+\frac{\beta}{K}\sum_{k=0}^{K-1}\mathbb{E}[\eta_{c}^{k}]^{2}$ (74)
@@ F.0.2 Rate of Constraint Violation
$0\leq\frac{1}{K}\left|\lambda_{K}-\frac{2}{\delta}\right|^{2}\leq\frac{1}{K}\left|{{\lambda_{0}-\frac{2}{\delta}}}\right|^{2}-\frac{2\beta}{K}\sum_{k=0}^{K-1}J_{c}(\theta_{k})\left(\lambda_{k}-\frac{2}{\delta}\right)+\beta^{2}$ (78)
@@ Appendix G Some Auxiliary Lemmas for the Proofs
[41, Lemma 14] For any ergodic MDP with mixing time $\tau_{\mathrm{mix}}$, the following holds $\forall(s,a)\in\mathcal{S}\times\mathcal{A}$, any policy $\pi$ and $\forall g\in\{r,c\}$.
@@
Applying Lemma G.1(b)(b) and Assumption 4.4(a)(a), we get
@@
[17, Lemma 3] We restate the problem (2) below for convenience.
@@
Notice that in Eq. (12), the dual update is projected onto the set $0,\frac{2}{\delta}]$ because the optimal dual variable for the parameterized problem is bounded in Lemma G.4. We note that our dual updating technique remains the same as [10], however we were able to achieve an improvement of global convergence rate due to the use of natural policy gradient and a more prudent choice of stepsizes ($\alpha,\beta$).